OpenStudy (anonymous):

Which is a power of 10? 1.001 0.210 1,000 5005

OpenStudy (unklerhaukus):

some powers of ten \[10^{-1}=0.1\]\[10^0=1\]\[10^1=10\]\[10^2=100\]

Directrix (directrix):

I guess the question is: Which of these is an integral power of 10. Following @UnkleRhaukus ' pattern 10³ = 1000

Directrix (directrix):

@KCB

OpenStudy (anonymous):

Thank you :) i was tryna help my sister i dont remember this stuff.

Directrix (directrix):

We are glad to help you both.